OpenStudy (anonymous):

Solve the equation 4q=48

OpenStudy (anonymous):

divide both sides by 4, and can you do it?

OpenStudy (anonymous):

q = 12

OpenStudy (saifoo.khan):

\[q = \frac{48}{4}\] \[q = 12\]

OpenStudy (karatechopper):

4q=48 4q/4=48/4 q=12